OpenStudy (anonymous):

log(6⋅11) Steps in solving this?

OpenStudy (anonymous):

use log (m.n) = log m+ log n

OpenStudy (anonymous):

put m= 6 and n= 11

OpenStudy (anonymous):

I suppose your teacher is looking for you to split it up this way log(6*11) = log(6) + log(11) = you'll have to use a calculator or tables at this point. Kind of a dumb question, though, as it's probably easier to do log(6*11) = log(66) = you'll have to use a calculator or tables at this point.

OpenStudy (anonymous):

thanks guys

OpenStudy (agent0smith):

What does the actual question say?